### v1

# BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis

Abstract
The widespread availability of code-mixed data can provide valuable insights into low-resource languages like Bengali, which have limited datasets. Sentiment analysis has been a fundamental text classification task across several languages for code-mixed data. However, there has yet to be a large-scale and diverse sentiment analysis dataset on code-mixed Bengali. We address this limitation by introducing BnSentMix, a sentiment analysis dataset on code-mixed Bengali consisting of 20,000 samples with $4$ sentiment labels from Facebook, YouTube, and e-commerce sites. We ensure diversity in data sources to replicate realistic code-mixed scenarios. Additionally, we propose $14$ baseline methods including novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving an overall accuracy of $69.8\%$ and an F1 score of $69.1\%$ on sentiment classification tasks. Detailed analyses reveal variations in performance across different sentiment labels and text types, highlighting areas for future improvement.
BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis
Sadia Alam, Md Farhan Ishmam, Navid Hasin Alvee, Md Shahnewaz Siddique, Md Azam Hossain, Abu Raihan Mostofa Kamal Network and Data Analysis Group, Department of Computer Science and Engineering, Islamic University of Technology, Dhaka, Bangladesh {sadiaalam,farhanishmam,navidhasin,shahnewaz,azam,raihan.kamal}@iut-dhaka.edu
1 Introduction
Figure 1: Examples of the four sentiment labels from our code-mixed Bengali-English dataset BnSentMix and the corresponding English translations. Red represents English words, blue represents Bengali words written in English alphabets, and cyan represents implicit words in the code-mixed text.
Dataset #Samples #SL #DS Filtering #Baselines PA
Hindi Joshi et al. (2016) 3.9k 3 1 Manual 10
Bengali Mandal et al. (2018) 5k 3 1 Manual 5
Tamil Chakravarthi et al. (2020b) 15.7k 5 1 langdetect 10
Malayalam Chakravarthi et al. (2020a) 6.7k 5 1 langdetect 10
Persian Sabri et al. (2021) 3.6k 3 1 Keywords search 3
Swiss Pustulka-Hunt et al. (2018) 963 3 1 Manual 7
BnSentMix (Ours) 20k 4 3 mBERT 14
Table 1: Comparison of the number of samples, #SL: Sentiment Labels, #DS: Data Sources, filtering method, number of baselines, and PA: Public Availability of various code-mixed (with English) sentiment analysis datasets.
In the rapidly evolving digital landscape, code-mixing has become increasingly prevalent, particularly in multilingual societies. Code-mixing is the phenomenon of alternating between two or more languages within a single conversation or sentence Thara and Poornachandran (2018). Code-mixing can occur in various forms, including intra-sentential switching, where words from different languages appear within the same sentence, and intra-word switching, where elements from other languages combine to form a single word Stefanich et al. (2019); Litcofsky and Van Hell (2017). Intra-sentential switching is more frequently observed in colloquial settings. One significant yet understudied domain of code-switching is Bengali-English code-mixed text.
We consider Fig. 1 where the sentences are examples of Bengali-English intra-sentential switching. Intra-word switching is observed in the negative sentiment example. Here, Movietar is considered a single word, whereas the Bengali sub-word tar indicates possession. We also observe several words in the transliterated text that are not explicitly written in the code-mixed text. These implicitly defined words increase the challenges in processing code-mixed Bengali-English texts.
With over 250 million native speakers globally, Bengali is the seventh most spoken language in the world but remains a low-resource language in terms of research. While typing texts, Bengali speakers often use Bengali-English code-mixed terms to express their thoughts in writing. Despite the prevalence of code-mixed text on social media platforms, e-commerce sites, and other digital spaces, there remains a notable scarcity of resources to analyze and process such data.
Sentiment analysis, the computational study of people’s opinions, sentiments, emotions, and attitudes expressed in written language, plays a critical role in various applications, including social media monitoring, customer feedback, market research, and public opinion analysis Wankhade et al. (2022). While substantial progress has been made in monolingual sentiment analysis Medhat et al. (2014); Birjali et al. (2021), the complexities introduced by code-mixed texts present unique challenges that current models struggle to address Barman et al. (2014). This is particularly true for Bengali-English code-mixed texts Chanda et al. (2016), which have not received adequate attention in existing research.
Table 1 highlights the limitations of Bengali-English code-mixed sentiment analysis datasets compared to other Indic-English code-mixed datasets. The only available Bengali dataset Mandal et al. (2018) is limited to $5k$ samples, $3$ sentiment labels, a single data source, $5$ baselines, and is not publicly available. The existing language detection tools also have severe limitations in filtering code-mixed Bengali-English. Tools like langdetect [^1] https://pypi.org/project/langdetect/ and Bengali phonetic parser [^2] https://github.com/porimol/bnbphoneticparser designed for general language identification and code-mixed Bengali identification struggled with the spelling nuances of code-mixed text. The limitations necessitated novel code-mixed Bengali detection tools.
Addressing these challenges, we present a novel and extensive Bengali-English code-mixed dataset designed for sentiment analysis. Our dataset is distinguished for its diversity, volume, and public availability. Data has been curated from YouTube, Facebook, and e-commerce platforms to encapsulate a broad spectrum of contexts and topics. The availability of a robust and diverse dataset is a critical step towards developing more effective and inclusive NLP tools, ultimately contributing to the better understanding and processing of code-mixed languages. Our contributions can be summarized as:
  • We present the first large-scale Bengali-English code-mixed dataset with 20,000 samples and four sentiment labels. The dataset will be publicly available to facilitate further research.
  • We describe an automated Bengali-English code-mixed text detection method using a pre-trained language model for efficient data filtering.
  • We establish 14 baselines including 3 novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving state-of-the-art performance on our dataset.
Figure 2: Dataset creation pipeline of the BnSentMix dataset.
2 Related Work
2.1 Code-Mixed Tasks and Methods
Code-mixed data can be the source of several text classification tasks Thara and Poornachandran (2018) with sentiment analysis Mahadzir et al. (2021) being one of the most popular ones. Other natural language processing tasks (NLP) on code-mixed data include hate speech detection Sreelakshmi et al. (2020), translation Gautam et al. (2021), part of speech tagging Vyas et al. (2014), emotion classification Ameer et al. (2022), language identification Mandal and Singh (2018), and speech synthesis Sitaram and Black (2016). Researchers also incorporated several techniques in processing code-mixed data including training data augmentation Gupta et al. (2021); Rizvi et al. (2021) and word embeddings specific to code-mixed text Pratapa et al. (2018).
2.2 Sentiment Analysis
The significance of sentiment analysis has grown substantially with the rise of social media, prompting extensive research on monolingual corpora. Studies explored various languages, including English Hu and Liu (2004); Wiebe et al. (2005); Jiang et al. (2019), Russian Rogers et al. (2018), German Cieliebak et al. (2017), Norwegian Mæhlum et al. (2019), and several Indian languages Agrawal and Awekar (2018); Rani et al. (2020). Multilingual sentiment analysis Dashtipour et al. (2016); Pustulka-Hunt et al. (2018) gained popularity with the recent advancements in multilingual language models Devlin et al. (2019); Conneau et al. (2020). Over the years, a vast number of code-mixing datasets with English have been proposed, as enlisted in Table 1.
2.3 Code-Mixing in Bengali
Bengali is often code-mixed with English Chanda et al. (2016) and Hindi Raihan et al. (2023). In Bengali-English code-mixing, English tokens are commonly used alongside romanized or transliterated Bengali Shibli et al. (2023). Sentiment analysis on code-mixed Bengali has limited studies, which either use small private datasets Mandal et al. (2018) or are performed in a multilingual setting Patra et al. (2018). Data augmentation techniques have also been explored to enhance code-mixed sentiment analysis datasets in Bengali Tareq et al. (2023). Emotion detection, a task similar to sentiment analysis, has also been studied in the context of code-mixed Bengali Raihan et al. (2024).
3 BnSentMix Dataset
3.1 Dataset Design
We designed BnSentMix, a large-scale code-mixed Bengali-English dataset with four sentiment labels. First, we ensured that the raw data was collected from multiple data sources, reflecting realistic code-mixed texts commonly found in digital spaces. Second, we labeled the data using four distinct sentiments: the commonly used positive, negative, and neutral sentiments, as well as a mixed sentiment. As illustrated in Fig. 1, the mixed sentiment represents instances where both positive and negative sentiments are conveyed within different parts of the text. We decided to include the mixed label because the associated sentences are frequently observed in everyday texts and cannot be correctly classified under the traditional sentiment labels.
Figure 3: Composition of data sources of the BnSentMix dataset.
3.2 Data Sourcing
We collected extensive user-generated content from YouTube comments, Facebook comments, and e-commerce site reviews. These data sources were chosen for their high engagement rates and diverse linguistic input. YouTube comments were scraped using the YouTube API. We used Facepager [^3] https://github.com/strohne/Facepager to extract comments from public Facebook posts, pages, and groups. Selenium [^4] https://selenium-python.readthedocs.io/ was employed to mimic human browsing behavior on e-commerce sites to scrape product reviews. We amassed over 3 million samples of user-generated content, forming the foundation for our dataset and subsequent analysis. Fig. 3 illustrates the composition of the aforementioned data sources.
3.3 Data Cleaning
We discard samples with four words or less and samples containing external URLs. Redundant whitespaces, special characters, and non-ASCII characters including emojis and emoticons are also removed. Consequent sequences of punctuation symbols are reduced to single instances. The English words are downcased unless they appear at the beginning of the sentence. However, we did not correct any form of typing or grammatical errors in our dataset to ensure the trained model is robust for practical scenarios. The data cleaning procedure excluding the preliminary filtering is described in Algorithm 1.
Algorithm 1 Clean Text
0:  $text\leftarrow$ Input text
0:  $text\leftarrow$ Preprocessed text
1:  $text\leftarrow text.lower()$Convert to lowercase}
2:  $text\leftarrow$ Remove all special characters except "?", ",", "!", and "."
3:  $text\leftarrow$ Reduce consecutive sequences of punctuations to a single instance
4:  $text\leftarrow$ Remove all non-ASCII characters
5:  $text\leftarrow$ Remove extra white spaces
6:  $text\leftarrow$ Capitalize the first letter after each period (.)
7:  return  $text$
3.4 Data Filtering
We construct a novel Bengali-English code-mix detection dataset and fine-tune pre-trained language models to filter code-mixed Bengali-English from the data sources. Pre-trained language models excel in text classification tasks and can address the shortcomings of current rule-based language detection tools.
3.4.1 Code-mix Detection Dataset
The fine-tuning dataset for code-mixed Bengali-English detection comprises $3$ data sources. We incorporate the Dakshina dataset Roark et al. (2020) which has a rich collection of Southeast Asian languages, including many Bengali-English code-mixed sentences. Secondly, we utilized a Kaggle English dataset [^5] https://www.kaggle.com/datasets/rtatman/english-word-frequency consisting of a wide range of English words and extended with a third source Mandal and Singh (2018). By integrating these diverse sources, we curated a comprehensive dataset of $100k$ words, ensuring a balanced mix of Bengali, English, and code-mixed Bengali-English text. To maintain the linguistic purity of code-mixed Bengali, we exclude sentences containing words that are not English or Bengali.
3.4.2 Code-mix Detection Results
We evaluate $3$ pre-trained models – the multilingual models, mBERT Devlin et al. (2019) and XLM-RoBERTa Conneau et al. (2020), and the Bengali-English model BanglishBERT Bhattacharjee et al. (2022). Table 2 reveals mBERT showing substantially higher accuracy and F1 score in code-mixed Bengali-English detection. We argue that the pre-trained multilingual capabilities of mBERT effectively handled the nuances of code-mixed Bengali-English text.
Model Acc(%) F1 Score
XLM-RoBERTa 89.60 0.8985
BanglishBERT 90.56 0.8961
mBERT 94.56 0.9403
Table 2: Comparison of the accuracy and F1 score of the code-mixed Bengali-English detection methods.
Algorithm 2 Detect Code-mixed Bengali
0:  $S\leftarrow$ List of sentences
0:  $model\leftarrow$ Pre-trained mBERT model
0:  $tokenizer\leftarrow$ Pre-trained mBERT tokenizer
0:  $pred\leftarrow$ Predicted class label (0 or 1)
1:  $b\_count\leftarrow 0$
2:  $w\_count\leftarrow 0$
3:  for each $sent$ in $S$ do
4:     $words\leftarrow$ split($sent$)
5:     for each $w$ in $words$ do
6:        $w\leftarrow$ preprocess($w$)
7:        if $w$ is empty then
8:           continue
9:        end if
10:        $w\_count\leftarrow w\_count+1$
11:        $inputs\leftarrow$ tokenize($w$)
12:        $outputs\leftarrow$ model($inputs$)
13:        $pred\_class\leftarrow$ argmax($outputs$)
14:        if $pred\_class==1$ then
15:           $b\_count\leftarrow b\_count+1$
16:        end if
17:     end for
18:  end for
19:  if $w\_count<4$ then
20:     return  0
21:  end if
22:  $b\_percent\leftarrow b\_count/w\_count$
23:  if $b\_percent\geq 0.3$ then
24:     return  1
25:  else
26:     return  0
27:  end if
3.4.3 Filtering Pipeline
Following Fig. 2, the code-mixed Bengali-English is cleaned and passed to the data filtering classifier. The classifier detects the code-mixed Bengali samples from the multilingual data source following Algorithm 2. The filtered code-mixed Bengali is manually verified to ensure the highest quality of data. To maintain linguistic purity, we exclude sentences containing words that are not English or Bengali. We discard samples containing hate speech and offensive content. We also discarded samples with a small proportion of code-mixing. Sentences are considered to be code-mixed Bengali-English if they contain at least 30% English words. We obtain 21,587 data samples after data filtering.
Statistic Value
Mean Character Length 62.77
Max Character Length 1985
Min Character Length 14
Mean Word Count 11.65
Max Word Count 368
Min Word Count 4
Unique Word Count 37734
Unique Sentence Count 21873
Table 3: Key statistics of the BnSentMix dataset.
3.5 Data Annotation
Each sample in our dataset has been annotated twice by two different annotators to ensure generalized sentiment is conveyed. To perform data annotation, we recruited $64$ annotators who had been provided hourly monetary compensation. The data annotators have at least a high-school degree (equivalent to Grade 12 education) and are familiar with social media and digital spaces. The annotators were asked to re-label the same $250$ samples. We measured the inter-annotator agreement level and found Cohen’s Kappa $\kappa=0.86$, indicating substantial agreement.
Figure 4: Distribution of sentiment labels in the BnSentMix dataset.
3.6 Annotation Validation
The annotated data has been further verified by a third annotator. In cases where the two independent annotations did not match, the third annotator would break the tie. This additional step ensured that any contentious samples were thoroughly reviewed and the final annotation reflected a consensus among the annotators. Samples labeled unfit for sentiment analysis after data validation have been discarded.
3.7 Dataset Statistics
After a rigorous evaluation process, 20,000 samples were finalized from the 21,587 samples, ensuring high-quality annotations for subsequent analysis and model training. Fig. 4 visualizes the label composition of the annotated dataset. An overview of the key statistics of the annotated dataset is shown in table-3.
3.8 Train-Validation-Test Set Split
To evaluate the performance of our models effectively, we split the dataset into $[70:15:15]$ training, validation, and test splits i.e. 14,000, 3,000, and 3,000 samples respectively.
4 Methodology and Experimental Setup
4.1 Baseline Models
We evaluate $14$ baselines encompassing traditional machine learning models, recurrent neural network variants, and transformer-based pre-trained language models, observed in table 4.
4.2 Fine-tuning Transformer Baselines
While the non-transformer models were trained from scratch, the transformer models leveraged pre-trained weights that were fine-tuned on our dataset. We perform task-specific fine-tuning Howard and Ruder (2018) by initializing the model parameters $\theta_{0}=\theta_{\text{pre}}$, and updating them following the gradient descent equation for $t+1$ iteration:
$\theta_{t+1}=\theta_{t}-\alpha\nabla_{\theta}J(\theta)$ (1)
where $\theta_{i}$ represents the parameters for the $i$ iteration, $\alpha$ is the learning rate and $J(\theta)$ is the derivative of the loss function with respect to the given parameters.
4.3 Further Pre-trained Transformer Models
We enhance the performance of the transformer baselines by incorporating Masked Language Modeling loss Devlin et al. (2019) to further pre-train the models on the unlabeled code-mixed data corpus scraped in sec-3.2. The MLM loss following dynamic masking can be formulated as:
$L_{\text{MLM}}=-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid\mathbf{x}_{\text{mask}})$ (2)
where, $x_{i}$ is the token at position $i$ masked in the input sequence, $\mathbf{x}_{\text{mask}}$ denotes the input sequence with dynamically masked tokens, $N$ is the total number of masked tokens. We ensure that samples from the test and validation sets are excluded from the further pre-training corpus. The further pre-trained models are referred to as Code-Mixed Bengali (CMB) transformer models and they follow the naming convention $x$-CMB, where $x$ represents a pre-trained transformer variant.
4.4 Evaluation Metrics
We use classification accuracy and F1-score for model evaluation – both well-known metrics for text classification Hossin and Sulaiman (2015).
4.4.1 Accuracy
Classification accuracy or accuracy, in short, quantifies the proportion of correctly classified predictions made by the classifier to the total predictions. The metric can be formulated:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (3)
where, TP, TN, FP, and FN represent the counts of True Positives, True Negatives, False Positives, and False Negatives respectively.
4.4.2 F1 Score
Accuracy can be misleading for imbalanced datasets. Instead, the F1 Score can provide a better estimate, computed as the harmonic mean of precision and recall.
[…]
where,
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (5)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (6)
4.5 Implementation Details
The models were trained on NVIDIA Tesla P100 GPUs with 16GB of memory. We followed the Huggingface implementation Wolf et al. (2019) for the pre-trained language models. All the models utilized Adam Optimizer Kingma and Ba (2014) with a training batch size of $32$.
The training configuration used most of the default hyperparameters. Logistic Regression, RNN, and LSTM models used the learning rate of $1E{-5}$ while the BERT-family language models used the learning rate of $1.5E{-6}$. The training time for each epoch varied from $8$ to $13$ minutes.
Model Validation Test
Acc Precision Recall F1 Acc Precision Recall F1
Machine Learning Models
Logistic Regression 0.668 0.656 0.668 0.662 0.667 0.614 0.667 0.639
Random Forest 0.672 0.661 0.672 0.666 0.648 0.635 0.648 0.641
SVM 0.694 0.676 0.694 0.685 0.660 0.637 0.660 0.648
Recurrent Neural Network Variants
RNN 0.406 0.308 0.406 0.350 0.401 0.352 0.401 0.375
LSTM 0.678 0.670 0.678 0.674 0.670 0.657 0.670 0.663
Pre-trained Transformer Models
XLM-RoBERTa 0.726 0.709 0.726 0.717 0.698 0.642 0.698 0.669
mBERT 0.726 0.713 0.726 0.719 0.694 0.675 0.694 0.684
BanglaBERT 0.721 0.668 0.721 0.693 0.698 0.642 0.698 0.669
BanglishBERT 0.694 0.715 0.694 0.704 0.686 0.653 0.686 0.669
DistilBERT 0.701 0.694 0.701 0.697 0.672 0.665 0.672 0.668
BERT 0.727 0.710 0.724 0.717 0.695 0.683 0.694 0.688
Code-Mixed Bengali (CMB) Transformer Models
mBERT-CMB 0.727 0.713 0.727 0.720 0.694 0.675 0.695 0.685
BanglaBERT-CMB 0.723 0.669 0.724 0.695 0.699 0.643 0.699 0.670
BERT-CMB 0.730 0.712 0.728 0.720 0.698 0.685 0.698 0.691
Table 4: Performance of the proposed baselines based on accuracy, precision, recall, and F1 score.
5 Results and Analysis
5.1 Performance Evaluation
Table 4 highlights the performance of the $14$ baselines with BERT-CMB being the best-performing model in terms of both accuracy and F1 score. We now analyze the category-wise model performance.
5.1.1 Machine Learning (ML) Models
The ML models provide simple baselines and achieve considerably high accuracy, with the Support Vector Machine (SVM) Vapnik (1995) achieving accuracy and F1 score on par with larger transformer-based models like BanglishBERT. The other two ML baselines Logistic Regression Cox (1958) and Random Forest Breiman (2001) achieve satisfactory performance with relatively simpler architectures. These ML baselines can be effective in resource-constrained scenarios.
5.1.2 Recurrent Neural Networks (RNNs)
RNN Hopfield (1982) underperformed compared to the other baselines. On the contrary, the performance of Long Short-Term Memory (LSTM) models Hochreiter and Schmidhuber (1997) was significantly higher in terms of both accuracy and F1 score. We argue that the long-term textual dependencies and the impact of vanishing and exploding gradients limited the performance of the RNN models.
5.1.3 Transformer-based Models
The best performance is achieved by the BERT model Devlin et al. (2019) pre-trained on an English corpus. The BERT model is closely followed by the multilingual models XLM-RoBERTa Conneau et al. (2020) and mBERT Devlin et al. (2019). We hypothesize that the low proportion of Bengali text in the multilingual pre-training corpus does not provide any significant advantage in code-mixed Bengali classification tasks.
In contrast, English pre-trained models like BERT exhibit a better understanding of the linguistic intricacies of English words used in code-mixed Bengali, thereby producing better performance than other multilingual and Bengali models. Similarly, the Bengali language models BanglaBERT Bhattacharjee et al. (2022) and BanglishBERT Bhattacharjee et al. (2022) are trained on Bengali and Bengali-English corpora respectively. Code-mixed Bengali uses English tokens and hence, the pre-training on Bengali tokens does not provide any significant advantage. The lighter version of BERT, DistilBERT Sanh et al. (2019) produces comparable but slightly worse results.
5.1.4 Code-Mixed Bengali Models
We selected the three best-performing transformer encoders and further pre-trained them following sec-4.3. The models were selected from each sub-category – mBERT as the best multilingual model, BanglaBERT as the best Bengali model, and BERT as the best English model. We find that further pre-training on the code-mixed data translates to an increase in performance for all three models. The only exception is the mBERT-CMB model showing insignificant improvement in accuracy and F1-score. Similar to the pre-trained baselines, the BERT-CMB is the best-performing model followed by mBERT-CMB and BanglaBERT-CMB. We argue that further pre-training in a larger code-mixed corpus might translate to better performance for multilingual models like mBERT or XLM-RoBERTa.
5.2 Training Loss Analysis
Figure 5 illustrates the training loss across $15$ epochs for the baselines. We observe that all models converge before reaching the $15^{th}$ epoch. The only exception is the LSTM model which shows a slight indication of being benefited by additional training epochs. Excluding DistilBERT, the other BERT family models converged relatively faster in the earlier epochs. For most models, training for $5$-$8$ epochs is appropriate to prevent overfitting.
Figure 5: Comparison of epoch-wise training loss of the established baselines.
5.3 Hyperparameter Tuning
The hyperparameters play a crucial role in optimizing the performance of deep learning models. We analyze the role of hyperparameters for the best-performing BERT-CMB model and optimize them to gain superior performance.
5.3.1 Fine-tuning Epochs
Figure 6 shows the effect of fine-tuning epochs on the performance metrics. Both metrics increase up to the $6^{th}$ and stabilize onwards, suggesting keeping the number of fine-tuning epochs between $5$ to $7$.
Figure 6: Epoch-wise training accuracy and F1 score curves of the best performing BERT-CMB model.
5.4 Learning Rate vs. Accuracy Analysis
Figure 7 illustrates the relationship between the learning rate and the accuracy for the BERT-CMB model. While tuning the hyperparameters, we observed the highest accuracy in the validation set at the learning rate of $1.25E{-6}$ which is slightly lower than the default learning rate of $1.5E{-6}$. The curve shows no particular advantage for higher or lower rates.
Figure 7: The effect of learning rate on the validation accuracy of the best performing BERT-CMB model.
6 Conclusion
We introduce BnSentMix, a novel sentiment analysis dataset tailored for code-mixed Bengali-English. Our work opens several potential research avenues for code-mixed Bengali. Researchers can explore other tasks, such as hate speech, offensive language, and abusive content detection on code-mixed data. Furthermore, the role of large language models and model ensembling techniques in processing code-mixed data can be further investigated. Our work addresses a significant gap for low-resource languages and sets a new standard for sentiment analysis in code-mixed Bengali-English.
Data Availability
Our dataset will be publicly available under the Creative Commons Attribution 4.0 International (CC BY 4.0). Any form of private data or personal identification information has been removed from the dataset to prevent privacy violations. We have ensured that the redistribution of social media data is consistent with the policies of the corresponding platforms.
Limitations
The label distribution of BnSentMix dataset is slightly imbalanced with only 9.2% samples labeled as mixed sentiment which can affect the performance of the model in classifying mixed sentiments. We also acknowledge that the sentiment of the annotator can be a source of bias during data annotation, though each data sample has been annotated twice by two different annotators, and annotation conflicts have been resolved by a data validator. Emojis and emoticons can convey sentiment but have been removed from our dataset. Their impact on the code-mixed sentiment can be explored in future works. Finally, complex sentiments involving sarcasm and humor are usually classified as neutral but can be subjective depending on the context.
Ethical Statement
The hired data annotators were given monetary compensation which was significantly higher than the minimum wage of the region. Each annotator was given only $630$ data samples to annotate with no time restrictions. This ensured that the annotator did not overwork during data annotation. Annotator sentiment is subject to long working hours and can affect the sentiment labeling. To prevent this, we mandated five-minute breaks after every twenty-minute interval and provided refreshments upon request.

### v3

@@
 Abstract
-The widespread availability of code-mixed data can provide valuable insights into low-resource languages like Bengali, which have limited datasets.
-Sentiment analysis has been a fundamental text classification task across several languages for code-mixed data.
-However, there has yet to be a large-scale and diverse sentiment analysis dataset on code-mixed Bengali.
-We address this limitation by introducing BnSentMix, a sentiment analysis dataset on code-mixed Bengali consisting of 20,000 samples with $4$ sentiment labels from Facebook, YouTube, and e-commerce sites.
-We ensure diversity in data sources to replicate realistic code-mixed scenarios.
-Additionally, we propose $14$ baseline methods including novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving an overall accuracy of $69.8\%$ and an F1 score of $69.1\%$ on sentiment classification tasks.
-Detailed analyses reveal variations in performance across different sentiment labels and text types, highlighting areas for future improvement.
-BnSentMix: A Diverse Bengali-English Code-Mixed Dataset for Sentiment Analysis
-Sadia Alam, Md Farhan Ishmam, Navid Hasin Alvee, Md Shahnewaz Siddique, Md Azam Hossain, Abu Raihan Mostofa Kamal Network and Data Analysis Group, Department of Computer Science and Engineering, Islamic University of Technology, Dhaka, Bangladesh {sadiaalam,farhanishmam,navidhasin,shahnewaz,azam,raihan.kamal}@iut-dhaka.edu
+The widespread availability of code-mixed data in digital spaces can provide valuable insights into low-resource languages like Bengali, which have limited annotated corpora.
+Sentiment analysis, a pivotal text classification task, has been explored across multiple languages, yet code-mixed Bengali remains underrepresented with no large-scale, diverse benchmark.
+Code-mixed text is particularly challenging as it requires the understanding of multiple languages and their interaction in the same text.
+We address this limitation by introducing BnSentMix, a sentiment analysis dataset on code-mixed Bengali comprising 20,000 samples with $4$ sentiment labels, sourced from Facebook, YouTube, and e-commerce sites.
+By aggregating multiple sources, we ensure linguistic diversity reflecting realistic code-mixed scenarios.
+We implement a novel automated text filtering pipeline using fine-tuned language models to detect code-mixed samples and expand code-mixed text corpora.
+We further propose baselines using machine learning, neural networks, and transformer-based language models.
+The availability of a diverse dataset is a critical step towards democratizing NLP and ultimately contributing to a better understanding of code-mixed languages.
+BnSentMix: A Diverse Bengali-English Code-Mixed Dataset
+for Sentiment Analysis
+Sadia Alam, Md Farhan Ishmam, Navid Hasin Alvee, Md Shahnewaz Siddique, Abu Raihan Mostofa Kamal, Md Azam Hossain Department of Computer Science and Engineering, Islamic University of Technology {sadiaalam,farhanishmam,navidhasin,shahnewaz,raihan.kamal,azam}@iut-dhaka.edu
 1 Introduction
@@
 Tools like langdetect [^1] https://pypi.org/project/langdetect/ and Bengali phonetic parser [^2] https://github.com/porimol/bnbphoneticparser designed for general language identification and code-mixed Bengali identification struggled with the spelling nuances of code-mixed text.
-The limitations necessitated novel code-mixed Bengali detection tools.
-Addressing these challenges, we present a novel and extensive Bengali-English code-mixed dataset designed for sentiment analysis.
-Our dataset is distinguished for its diversity, volume, and public availability.
+Addressing these challenges, our contribution can be summarized:
+  • We present, BnSentMix, a novel Bengali-English code-mixed dataset comprising 20,000 samples and 4 sentiment labels for sentiment analysis.
 Data has been curated from YouTube, Facebook, and e-commerce platforms to encapsulate a broad spectrum of contexts and topics.
-The availability of a robust and diverse dataset is a critical step towards developing more effective and inclusive NLP tools, ultimately contributing to the better understanding and processing of code-mixed languages.
-Our contributions can be summarized as:
-  • We present the first large-scale Bengali-English code-mixed dataset with 20,000 samples and four sentiment labels. The dataset will be publicly available to facilitate further research.
-  • We describe an automated Bengali-English code-mixed text detection method using a pre-trained language model for efficient data filtering.
-  • We establish 14 baselines including 3 novel transformer encoders further pre-trained on code-mixed Bengali-English, achieving state-of-the-art performance on our dataset.
+  • Following the intricacies of code-mixed test, visualized in Fig.
+1, we propose a novel automated code-mixed text detection pipeline using fine-tuned language models, reaching an accuracy of 94.56%.
+  • We establish 11 baselines including classical machine learning, neural network, and pre-trained transformer-based models, with BERT achieving accuracy and F1 score of 69.5% and 68.8% respectively.
 Figure 2: Dataset creation pipeline of the BnSentMix dataset.
 2 Related Work
-2.1 Code-Mixed Tasks and Methods
+2.1 Code-Mixing
 Code-mixed data can be the source of several text classification tasks Thara and Poornachandran (2018) with sentiment analysis Mahadzir et al.
@@
 (2022), language identification Mandal and Singh (2018), and speech synthesis Sitaram and Black (2016).
-Researchers also incorporated several techniques in processing code-mixed data including training data augmentation Gupta et al.
+Researchers also incorporate training data augmentation Gupta et al.
 (2021); Rizvi et al.
-(2021) and word embeddings specific to code-mixed text Pratapa et al.
-(2018).
+(2021) and code-mix word embeddings Pratapa et al.
+(2018) to process code-mixed texts.
 2.2 Sentiment Analysis
-The significance of sentiment analysis has grown substantially with the rise of social media, prompting extensive research on monolingual corpora.
+The significance of sentiment analysis has grown with the rise of social media, prompting extensive research on monolingual corpora.
 Studies explored various languages, including English Hu and Liu (2004); Wiebe et al.
@@
 (2017), Norwegian Mæhlum et al.
-(2019), and several Indian languages Agrawal and Awekar (2018); Rani et al.
+(2019), several Indian languages Agrawal and Awekar (2018); Rani et al.
+(2020), and Bengali Fahim (2023); Kabir et al.
 (2023).
@@
 (2020).
-Over the years, a vast number of code-mixing datasets with English have been proposed, as enlisted in Table 1.
 2.3 Code-Mixing in Bengali
@@
 In Bengali-English code-mixing, English tokens are commonly used alongside romanized or transliterated Bengali Shibli et al.
+(2023); Fahim et al.
+(2024), which is often back-transliterated before processing Haider et al.
 (2024).
-Sentiment analysis on code-mixed Bengali has limited studies, which either use small private datasets Mandal et al.
-(2018) or are performed in a multilingual setting Patra et al.
+Sentiment analysis on code-mixed Bengali has limited studies, either using small private datasets Mandal et al.
+(2018) or performed in a multilingual setting Patra et al.
 (2018).
@@
 3 BnSentMix Dataset
-3.1 Dataset Design
-We designed BnSentMix, a large-scale code-mixed Bengali-English dataset with four sentiment labels.
-First, we ensured that the raw data was collected from multiple data sources, reflecting realistic code-mixed texts commonly found in digital spaces.
-Second, we labeled the data using four distinct sentiments: the commonly used positive, negative, and neutral sentiments, as well as a mixed sentiment.
+The BnSentMix data has been collected from multiple data sources to reflect realistic code-mixed texts commonly found in digital spaces.
+We labeled the data using four distinct sentiments: the commonly used positive, negative, and neutral sentiments, as well as a mixed sentiment.
 As illustrated in Fig.
@@
 However, we did not correct any form of typing or grammatical errors in our dataset to ensure the trained model is robust for practical scenarios.
-The data cleaning procedure excluding the preliminary filtering is described in Algorithm 1.
+The data cleaning procedure has been formally described in Algo.
+1.
 Algorithm 1 Clean Text
@@
 3.3 Data Filtering
-We construct a novel Bengali-English code-mix detection dataset and fine-tune pre-trained language models to filter code-mixed Bengali-English from the data sources.
-Pre-trained language models excel in text classification tasks and can address the shortcomings of current rule-based language detection tools.
-3.4.1 Code-mix Detection Dataset
-The fine-tuning dataset for code-mixed Bengali-English detection comprises $3$ data sources.
-We incorporate the Dakshina dataset Roark et al.
-(2020) which has a rich collection of Southeast Asian languages, including many Bengali-English code-mixed sentences.
-Secondly, we utilized a Kaggle English dataset [^5] https://www.kaggle.com/datasets/rtatman/english-word-frequency consisting of a wide range of English words and extended with a third source Mandal and Singh (2018).
-By integrating these diverse sources, we curated a comprehensive dataset of $100k$ words, ensuring a balanced mix of Bengali, English, and code-mixed Bengali-English text.
-To maintain the linguistic purity of code-mixed Bengali, we exclude sentences containing words that are not English or Bengali.
-3.4.2 Code-mix Detection Results
-We evaluate $3$ pre-trained models – the multilingual models, mBERT Devlin et al.
-(2019) and XLM-RoBERTa Conneau et al.
-(2020), and the Bengali-English model BanglishBERT Bhattacharjee et al.
-(2022).
-Table 2 reveals mBERT showing substantially higher accuracy and F1 score in code-mixed Bengali-English detection.
-We argue that the pre-trained multilingual capabilities of mBERT effectively handled the nuances of code-mixed Bengali-English text.
-Model Acc(%) F1 Score
-XLM-RoBERTa 89.60 0.8985
-BanglishBERT 90.56 0.8961
-mBERT 94.56 0.9403
-Table 2: Comparison of the accuracy and F1 score of the code-mixed Bengali-English detection methods.
+We construct a novel Bengali-English code-mix detection dataset and fine-tune pre-trained language models to automatically filter code-mixed Bengali-English.
+Detecting these texts can pose significant challenges: (i) rule-based methods struggle with intra-word switching (ii) romanized Bengali or English samples may be incorrectly classified as code-mixed text by automated methods, and (iii) samples from a third language often bypass the filtering process.
+Our approach addresses these challenges by incorporating pre-trained language models, which excel in intricate text detection settings.
+Algo.
+2 outlines the data filtering pipeline.
 Algorithm 2 Detect Code-mixed Bengali
@@
 27:  end if
-3.4.3 Filtering Pipeline
-Following Fig.
-2, the code-mixed Bengali-English is cleaned and passed to the data filtering classifier.
-The classifier detects the code-mixed Bengali samples from the multilingual data source following Algorithm 2.
-The filtered code-mixed Bengali is manually verified to ensure the highest quality of data.
-To maintain linguistic purity, we exclude sentences containing words that are not English or Bengali.
-We discard samples containing hate speech and offensive content.
-We also discarded samples with a small proportion of code-mixing.
-Sentences are considered to be code-mixed Bengali-English if they contain at least 30% English words.
-We obtain 21,587 data samples after data filtering.
+3.3.1 Code-mix Detection Dataset
+The fine-tuning dataset for code-mixed Bengali-English detection comprises $3$ data sources.
+We incorporate the Dakshina dataset Roark et al.
+(2020) which has a rich collection of Southeast Asian languages, including many Bengali-English code-mixed sentences.
+Secondly, we utilized a Kaggle English dataset [^5] https://www.kaggle.com/datasets/rtatman/english-word-frequency consisting of a wide range of English words and extended with a third source Mandal and Singh (2018).
+By integrating these diverse sources, we curated a comprehensive dataset of $100k$ words, ensuring a balanced mix of Bengali, English, and code-mixed Bengali-English words.
+To maintain the linguistic purity of code-mixed Bengali-English, we exclude sentences containing words that are neither English nor Bengali, e.g.
+Hindi words.
+3.3.2 Code-mix Detection Results
+We evaluate $3$ pre-trained models – the multilingual models, mBERT Devlin et al.
+(2019) and XLM-RoBERTa Conneau et al.
+(2020), and the Bengali-English model BanglishBERT Bhattacharjee et al.
+(2022).
+Table 2 reveals mBERT showing substantially higher accuracy and F1 score in code-mixed Bengali-English detection.
+We argue that the pre-trained multilingual capabilities of mBERT effectively handled the nuances of code-mixed Bengali-English text.
+Model Acc(%) F1 Score
+XLM-RoBERTa 89.60 0.8985
+BanglishBERT 90.56 0.8961
+mBERT 94.56 0.9403
+Table 2: Comparison of the accuracy and F1 score of the code-mixed Bengali-English detection methods.
+Figure 4: Distribution of sentiment labels in the BnSentMix dataset.
+3.4 Data Annotation
+Each sample in our dataset has been annotated twice by two different annotators to ensure generalized sentiment is conveyed.
+In cases where the two independent annotations did not match, a third annotator would break the tie.
+To perform data annotation, we recruited $64$ annotators who had been provided hourly monetary compensation.
+The data annotators have at least a high-school degree (equivalent to Grade 12 education) and are familiar with social media and digital spaces.
+The annotators were asked to re-label the same $250$ samples to measure inter-annotator agreement.
+We measured the agreement score using Cohen’s Kappa $\kappa=0.86$, indicating substantial agreement.
+3.5 Dataset Statistics
+Fig.
+4 visualizes the label composition of the annotated dataset.
+An overview of the key statistics of the annotated dataset is shown in table-3.
+We split the dataset into $[70:15:15]$ training, validation, and test splits i.e.
+14,000, 3,000, and 3,000 samples respectively.
+4 Methodology and Experimental Setup
+4.1 Baseline Models
+We evaluate $11$ baselines encompassing traditional machine learning models, recurrent neural network variants, and transformer-based pre-trained language models, observed in table 4.
+All the pre-trained models were fine-tuned on our dataset.
+4.2 Evaluation Metrics
+We use classification accuracy and F1-score for model evaluation – both well-known metrics for text classification Hossin and Sulaiman (2015).
 Statistic Value
@@
 Unique Word Count 37734
-Unique Sentence Count 21873
+Unique Sentence Count 20000
 Table 3: Key statistics of the BnSentMix dataset.
-3.5 Data Annotation
-Each sample in our dataset has been annotated twice by two different annotators to ensure generalized sentiment is conveyed.
-To perform data annotation, we recruited $64$ annotators who had been provided hourly monetary compensation.
-The data annotators have at least a high-school degree (equivalent to Grade 12 education) and are familiar with social media and digital spaces.
-The annotators were asked to re-label the same $250$ samples.
-We measured the inter-annotator agreement level and found Cohen’s Kappa $\kappa=0.86$, indicating substantial agreement.
-Figure 4: Distribution of sentiment labels in the BnSentMix dataset.
-3.6 Annotation Validation
-The annotated data has been further verified by a third annotator.
-In cases where the two independent annotations did not match, the third annotator would break the tie.
-This additional step ensured that any contentious samples were thoroughly reviewed and the final annotation reflected a consensus among the annotators.
-Samples labeled unfit for sentiment analysis after data validation have been discarded.
-3.7 Dataset Statistics
-After a rigorous evaluation process, 20,000 samples were finalized from the 21,587 samples, ensuring high-quality annotations for subsequent analysis and model training.
-Fig.
-4 visualizes the label composition of the annotated dataset.
-An overview of the key statistics of the annotated dataset is shown in table-3.
-3.8 Train-Validation-Test Set Split
-To evaluate the performance of our models effectively, we split the dataset into $[70:15:15]$ training, validation, and test splits i.e. 14,000, 3,000, and 3,000 samples respectively.
-4 Methodology and Experimental Setup
-4.1 Baseline Models
-We evaluate $14$ baselines encompassing traditional machine learning models, recurrent neural network variants, and transformer-based pre-trained language models, observed in table 4.
-4.2 Fine-tuning Transformer Baselines
-While the non-transformer models were trained from scratch, the transformer models leveraged pre-trained weights that were fine-tuned on our dataset.
-We perform task-specific fine-tuning Howard and Ruder (2018) by initializing the model parameters $\theta_{0}=\theta_{\text{pre}}$, and updating them following the gradient descent equation for $t+1$ iteration:
-$\theta_{t+1}=\theta_{t}-\alpha\nabla_{\theta}J(\theta)$ (1)
-where $\theta_{i}$ represents the parameters for the $i$ iteration, $\alpha$ is the learning rate and $J(\theta)$ is the derivative of the loss function with respect to the given parameters.
-4.3 Further Pre-trained Transformer Models
-We enhance the performance of the transformer baselines by incorporating Masked Language Modeling loss Devlin et al.
-(2019) to further pre-train the models on the unlabeled code-mixed data corpus scraped in sec-3.2.
-The MLM loss following dynamic masking can be formulated as:
-$L_{\text{MLM}}=-\frac{1}{N}\sum_{i=1}^{N}\log P(x_{i}\mid\mathbf{x}_{\text{mask}})$ (2)
-where, $x_{i}$ is the token at position $i$ masked in the input sequence, $\mathbf{x}_{\text{mask}}$ denotes the input sequence with dynamically masked tokens, $N$ is the total number of masked tokens.
-We ensure that samples from the test and validation sets are excluded from the further pre-training corpus.
-The further pre-trained models are referred to as Code-Mixed Bengali (CMB) transformer models and they follow the naming convention $x$-CMB, where $x$ represents a pre-trained transformer variant.
-4.4 Evaluation Metrics
-We use classification accuracy and F1-score for model evaluation – both well-known metrics for text classification Hossin and Sulaiman (2015).
-4.4.1 Accuracy
-Classification accuracy or accuracy, in short, quantifies the proportion of correctly classified predictions made by the classifier to the total predictions. The metric can be formulated:
-$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (3)
-where, TP, TN, FP, and FN represent the counts of True Positives, True Negatives, False Positives, and False Negatives respectively.
-4.4.2 F1 Score
-Accuracy can be misleading for imbalanced datasets. Instead, the F1 Score can provide a better estimate, computed as the harmonic mean of precision and recall.
-[…]
-where,
-$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (5)
-$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (6)
 4.3 Implementation Details
@@
        LSTM 0.678 0.670 0.678 0.674 0.670 0.657 0.670 0.663
-Pre-trained Transformer Models
+Multilingual Language Models
        XLM-RoBERTa 0.726 0.709 0.726 0.717 0.698 0.642 0.698 0.669
        mBERT 0.726 0.713 0.726 0.719 0.694 0.675 0.694 0.684
+Bangla Language Models
        BanglaBERT 0.721 0.668 0.721 0.693 0.698 0.642 0.698 0.669
        BanglishBERT 0.694 0.715 0.694 0.704 0.686 0.653 0.686 0.669
+English Language Models
        DistilBERT 0.701 0.694 0.701 0.697 0.672 0.665 0.672 0.668
        BERT 0.727 0.710 0.724 0.717 0.695 0.683 0.694 0.688
-Code-Mixed Bengali (CMB) Transformer Models
-mBERT-CMB 0.727 0.713 0.727 0.720 0.694 0.675 0.695 0.685
-BanglaBERT-CMB 0.723 0.669 0.724 0.695 0.699 0.643 0.699 0.670
-BERT-CMB 0.730 0.712 0.728 0.720 0.698 0.685 0.698 0.691
 Table 4: Performance of the proposed baselines based on accuracy, precision, recall, and F1 score.
@@
 5.1 Performance Evaluation
-Table 4 highlights the performance of the $14$ baselines with BERT-CMB being the best-performing model in terms of both accuracy and F1 score. We now analyze the category-wise model performance.
+Table 4 highlights the performance of the $11$ baselines with BERT achieving the best performance in terms of both accuracy and F1 score. We now analyze the category-wise model performance.
 5.1.1 Machine Learning (ML) Models
@@
 We argue that the long-term textual dependencies and the impact of vanishing and exploding gradients limited the performance of the RNN models.
+Figure 5: Comparison of epoch-wise training loss of the established baselines.
 5.1.3 Transformer-based Models
@@
 We hypothesize that the low proportion of Bengali text in the multilingual pre-training corpus does not provide any significant advantage in code-mixed Bengali classification tasks.
-In contrast, English pre-trained models like BERT exhibit a better understanding of the linguistic intricacies of English words used in code-mixed Bengali, thereby producing better performance than other multilingual and Bengali models.
+In contrast, English pre-trained models like BERT exhibit better understanding of the linguistic intricacies of English words used in code-mixed Bengali, thereby producing better performance than other multilingual and Bengali models.
 Similarly, the Bengali language models BanglaBERT Bhattacharjee et al.
@@
 (2019) produces comparable but slightly worse results.
-5.1.4 Code-Mixed Bengali Models
-We selected the three best-performing transformer encoders and further pre-trained them following sec-4.3.
-The models were selected from each sub-category – mBERT as the best multilingual model, BanglaBERT as the best Bengali model, and BERT as the best English model.
-We find that further pre-training on the code-mixed data translates to an increase in performance for all three models.
-The only exception is the mBERT-CMB model showing insignificant improvement in accuracy and F1-score.
-Similar to the pre-trained baselines, the BERT-CMB is the best-performing model followed by mBERT-CMB and BanglaBERT-CMB.
-We argue that further pre-training in a larger code-mixed corpus might translate to better performance for multilingual models like mBERT or XLM-RoBERTa.
 5.2 Training Loss Analysis
@@
 For most models, training for $5$-$8$ epochs is appropriate to prevent overfitting.
-Figure 5: Comparison of epoch-wise training loss of the established baselines.
-5.3 Hyperparameter Tuning
-The hyperparameters play a crucial role in optimizing the performance of deep learning models.
-We analyze the role of hyperparameters for the best-performing BERT-CMB model and optimize them to gain superior performance.
-5.3.1 Fine-tuning Epochs
-Figure 6 shows the effect of fine-tuning epochs on the performance metrics.
-Both metrics increase up to the $6^{th}$ and stabilize onwards, suggesting keeping the number of fine-tuning epochs between $5$ to $7$.
-Figure 6: Epoch-wise training accuracy and F1 score curves of the best performing BERT-CMB model.
-5.4 Learning Rate vs. Accuracy Analysis
-Figure 7 illustrates the relationship between the learning rate and the accuracy for the BERT-CMB model.
-While tuning the hyperparameters, we observed the highest accuracy in the validation set at the learning rate of $1.25E{-6}$ which is slightly lower than the default learning rate of $1.5E{-6}$.
-The curve shows no particular advantage for higher or lower rates.
-Figure 7: The effect of learning rate on the validation accuracy of the best performing BERT-CMB model.
 6 Conclusion
@@
 Researchers can explore other tasks, such as hate speech, offensive language, and abusive content detection on code-mixed data.
-Furthermore, the role of large language models and model ensembling techniques in processing code-mixed data can be further investigated.
 Our work addresses a significant gap for low-resource languages and sets a new standard for sentiment analysis in code-mixed Bengali-English.
@@
 The label distribution of BnSentMix dataset is slightly imbalanced with only 9.2% samples labeled as mixed sentiment which can affect the performance of the model in classifying mixed sentiments.
-We also acknowledge that the sentiment of the annotator can be a source of bias during data annotation, though each data sample has been annotated twice by two different annotators, and annotation conflicts have been resolved by a data validator.
-Emojis and emoticons can convey sentiment but have been removed from our dataset.
-Their impact on the code-mixed sentiment can be explored in future works.
-Finally, complex sentiments involving sarcasm and humor are usually classified as neutral but can be subjective depending on the context.
+We also acknowledge that the sentiment of the annotator can be a source of bias during data annotation, though each data sample has been annotated twice by two different annotators, and annotation conflicts have been resolved by a third annotator.
 Ethical Statement
-The hired data annotators were given monetary compensation which was significantly higher than the minimum wage of the region.
-Each annotator was given only $630$ data samples to annotate with no time restrictions.
+The hired data annotators were compensated significantly higher than the region’s minimum wage.
+Each annotator was only given around $630$ data samples with no time restrictions.
 This ensured that the annotator did not overwork during data annotation.
-Annotator sentiment is subject to long working hours and can affect the sentiment labeling.
+Annotator sentiment is subject to long working hours and can affect sentiment labeling.
 To prevent this, we mandated five-minute breaks after every twenty-minute interval and provided refreshments upon request.
+Acknowledgements
+Our work is supported by the Islamic University of Technology Research Seed Grants (IUT RSG) (Ref: REASP/IUT-RSG/2022/OL/07/012).
+We sincerely appreciate Mohammed Saidul Islam and Md Mezbaur Rahman for guidance and Nejd Khadija for proofreading our work.